OpenStudy (anonymous):

simplify square root 108 minus 3 square root 8 minus square root 27

OpenStudy (anonymous):

whats the answer bro

OpenStudy (anonymous):

\(\Large \sqrt{108}-3\sqrt8-\sqrt27=\sqrt{36\times3}-3\sqrt{4\times2}-\sqrt{9\times3}\) \(\Large =\sqrt{6^2} \sqrt3 -3\sqrt{2^2}\sqrt2-\sqrt{3^2}\sqrt3=6\sqrt3-6\sqrt2-3\sqrt3=3\sqrt3-6\sqrt2\)

OpenStudy (anonymous):

the answer is: \(\Large 3\sqrt3-6\sqrt2\)